# It’s High Time : A Survey of Temporal Information Retrieval and Question Answering

### v2

@@ Abstract
Time plays a critical role in how information is generated, retrieved, and interpreted. In this survey, we provide a comprehensive overview of Temporal Information Retrieval and Temporal Question Answering, two research areas aimed at handling and understanding time-sensitive information. As the amount of time-stamped content from sources like news articles, web archives, and knowledge bases increases, systems must address challenges such as detecting temporal intent, normalizing time expressions, ordering events, and reasoning over evolving or ambiguous facts. These challenges are critical across many dynamic and time-sensitive domains, from news and encyclopedias to science, history, and social media. We review both traditional approaches and modern neural methods, including those that use transformer models and Large Language Models (LLMs). We also review recent advances in temporal language modeling, multi-hop reasoning, and retrieval-augmented generation (RAG), alongside benchmark datasets and evaluation strategies that test temporal robustness, recency awareness, and generalization.
@@
: A Survey of Temporal Information Retrieval and Question Answering
@@ 1 Introduction
From analyzing centuries-old texts, understanding historical events, to answering questions about emerging developments, time shapes how we seek and interpret information. As digital content continues to grow exponentially across time-stamped sources like news archives, social media, and knowledge bases, the ability to process and reason over temporal information has become essential Alonso et al. (2007). Temporal IR, which searches time-stamped documents, and Temporal QA, which answers time-sensitive queries, together address these needs. Both, collectively referred as Temporal IR/QA, aim to incorporate time-awareness to adapt results to specific periods and resolve time-sensitive queries Campos et al. (2014).
Temporal IR/QA faces distinct challenges that set it apart from standard IR/QA settings. These include identifying temporal intent in queries, interpreting expressions such as "post-World War II" or "in 1998," and modeling relationships between events and their timelines Berberich et al. (2010). Queries may target past, present, or future events, and require systems to identify relevant time frames, order events, and resolve implicit temporal cues. Overcoming these obstacles demands methods that extend beyond traditional keyword-based search and basic retrieval techniques.
For Example, in Figure 1, Q1: "At what age did Obama win the Nobel Peace Prize?" requires constructing a chronology of events by identifying and grounding two temporal anchors, Obama’s birth year (1961) and the year he received the Nobel Peace Prize (2009). The model must then establish a temporal relationship and apply reasoning to compute the answer: 48 years old. Q2: “What does President Obama’s climate policy tell us about how the U.S. viewed climate change during his late years of service?” demands understanding of the query’s intended time and contextual temporal grounding. The model must recognize relative temporal expressions such as “today, next week” and associate them with a reference time such as the document’s publication date. It also needs to retrieve or reason over documents written during the relevant policy timeframe, reconstructing the contemporaneous narrative.
@@
Research in Temporal IR/QA has evolved significantly, building on early foundations to address increasingly complex temporal challenges. Initial efforts relied on rule-based systems Harabagiu and Bejan (2005) and statistical models Berberich et al. (2010) that used document timestamps and hand-crafted rules to interpret time-related information Li and Croft (2003). While these methods established key principles, they struggled to scale or handle diverse temporal contexts. The rise of pre-trained language models has transformed the field by enabling robust temporal reasoning Jain et al. (2023), event sequencing Lin et al. (2021), and adaptation to evolving knowledge Han et al. (2021). These advancements paved the way for more dynamic and scalable temporal systems.
While prior surveys111For a discussion of previous related surveys, we refer the reader to Section A.1. have explored general IR/QA methods Robertson et al. (2009); Xiong et al. (2020); Formal et al. (2021), or focused narrowly on specific aspects of temporal processing within one of these fields Kobayashi and Takeda (2000); Kanhabua et al. (2015); Campos et al. (2014), a comprehensive and unified overview of Temporal IR/QA is long overdue. The most recent dedicated survey in this area was published nearly a decade ago Campos et al. (2014). Since then, the field has grown substantially, driven by advances in Language Models, new datasets, and complex temporal tasks. A survey that captures recent advancements and outlines future research directions is then essential to foster progress and guide the community in developing time-aware systems. Our paper addresses this gap. We trace the evolution from traditional to neural approaches, highlight advances in tasks such as event dating, temporal modeling, and knowledge updating, and outline emerging challenges. In Figure 2, we portray a taxonomy overviewing temporal tasks, datasets, and approaches we will discuss in our review.
@@ 2 Key Concepts
We first introduce the core concepts related to Temporal IR/QA.
@@
Temporal Information Retrieval (TIR) aims to retrieve documents that are not only topically relevant but also aligned with the query’s temporal intent. Temporal intent may be explicit such as “Olympics 2024”, or implicit ones, such as “latest Apple earnings”. TIR relies on different temporal signals such as document timestamps (publication dates), temporal expressions (“March 2023”), and event mentions (“2024 Olympics”) to assess a document’s temporal relevance indicating how well its temporal scope matches the query Kanhabua and Nørvåg (2008); Singh et al. (2016).
@@
Temporal Question Answering (TQA) focuses on answering questions with temporal constraints, either explicitly stated, such as "Who won the Nobel Prize in Physics in 2020?" or implied, for instance, "What are the latest US climate policies?". Success in TQA requires understanding the question’s temporal intent and retrieving documents relevant to the corresponding time frame or ones published around that time.
@@
Temporal IR/QA rely on diverse temporal elements. Temporal signals are, in general, defined as features that convey time-related information in text. These include explicit temporal expressions like "March 2023" (used for indexing and filtering), implicit cues such as "recently" (requiring contextual interpretation), relative expressions like "last week" (necessitating to be anchored to a reference point), event-based references such as "2024 Olympics" (linking to known event timelines), and temporal metadata such as document timestamps which indicate publication time and often serve as proxies for judging freshness of content.
The concept of document focus time Jatowt et al. (2013) is crucial here. It denotes the specific time point or interval a document relates to. For example, a 2013’s publication discussing the 2010 Academy Awards has a focus time of 2010 while having a timestamp of 2013. Accurate focus time estimation of documents, using techniques like burst detection, temporal expression analysis, or timestamping named entities, enhances answer precision, especially in news or historical corpora Wang et al. (2020).
forked edges,
for tree=
grow=east,
reversed=true,
anchor=base west,
parent anchor=east,
child anchor=west,
base=left,
font=,
rectangle,
rounded corners,
align=left,
minimum width=0.1em,
edge+=darkgray, line width=0.8pt,
s sep=2pt,
inner xsep=2pt,
inner ysep=2pt,
draw=hidden-draw,
where level=1text width=4.7em,font=,,
where level=2text width=5em,font=,,
where level=3text width=6.8em,font=,,
where level=4text width=10em,font=,,
where level=5text width=6.8em,font=,,
Temporal IR/QA, draw=black, fill=gray!10, thick, text=black, ver
Dataset &
Benchmarks ( §3),
color=orange!90, fill=orange!20, thick, text=black
Temporal Document
Collections ( §3.1),
color=orange!90, fill=orange!20, thick, text=black
TQA Datasets ( §3.2),
color=orange!90, fill=orange!20, thick, text=black
TIR Dataset ( §3.3),
color=orange!90, fill=orange!20, thick, text=black
Temporal
Tasks ( §4),
color=cyan!100, fill=cyan!15, thick, text=black
Prediction,
color=cyan!100, fill=cyan!15, thick, text=black
Approaches ( §5),
color=red!80, fill=red!15, thick, text=black
Rule-Based &
Statistical Models ( §5.1),
color=red!80, fill=red!15, thick, text=black
Temporal Language
Models ( §5.2),
color=red!80, fill=red!15, thick, text=black
Temporal RAG ( §5.2.1),
color=red!80, fill=red!15, thick, text=black
Temporal Reasoning
( §5.2.2),
color=red!80, fill=red!15, thick, text=black
@@
Dataset #Questions Knowledge Source Creation Method Answer Type Time Frame Temporal Metadata Multi-Hop NewsQA Trischler et al. (2017) 119k News CS Freeform 2007-2015 ✗ ✗ TDDiscourse Naik et al. (2019) 6.1k News CS Extractive Unspecified ✗ ✗ TORQUE Ning et al. (2020) 21k News CS Abstractive - ✗ ✗ ArchivalQA Wang et al. (2022) 532k News AG Extractive 1987-2007 ✓ ✗ TimeQA Chen et al. (2021) 41.2K Wikipedia AG Extractive 1367-2018 ✗ ✗ TiQ Jia et al. (2024) 10K Wikipedia AG Freebase Unspecified ✗ ✗ TempQuestions Jia et al. (2018) 1.2k Freebase AG Extractive Unspecified ✗ ✓ TemporalQuestions Wang et al. (2021a) 1K News CS Extrcative 1987-2007 ✓ ✗ TempLAMA Dhingra et al. (2022) 50k News CS Extractive 2010-2020 ✓ ✗ ComplexTempQA Gruber et al. (2024) 100,228k Wikipedia AG Extractive 1987-2023 ✓ ✓ MenatQA Wei et al. (2023) 2.8k Wikipedia AG Extractive 1367-2018 ✗ ✗ PAT-Question Meem et al. (2024) 6,1k Wikipedia CS Extractive - ✗ ✓ TempTabQA Gupta et al. (2023) 11.4k Wikipedia Info box CS Abstractive - ✗ ✗ SituatedQA Zhang and Choi (2021) 12.2k Wikipedia CS – ✗ ✗ UnSeenTimeQA Uddin et al. (2024) 3.6k Synthetic AG Abstractive - ✗ ✓ ChroniclingAmericaQA Piryani et al. (2024b) 485k News AG Extractive 1800-1920 ✓ ✗ FRESHQA Vu et al. (2024) 600 Google Search CS - - ✗ ✓ COTEMPQA Su et al. (2024) 4.7k Wikidata CS Abstractive ✗ ✓ Test of Time (ToT) Fatemi et al. (2024) 1.8k Synthetic AG Abstractive - ✗ ✓ TIMEDAIL Qin et al. (2021) 1.1k DailyDialog CS Multiple-choice - ✗ ✗ Complex-TR Tan et al. (2024) 10.8 Wikipedia+Google Search AG Multi-answer ✗ ✓ StreamingQA Liska et al. (2022) 147k News CS Extractive 2007-2020 ✓ ✓ TRACIE Zhou et al. (2021) 5.4k Wikipedia CS abstractive ✗ ✗ ForecastQA Jin et al. (2021) 10.3k News CS Multiple-Choice 2015-2019 ✓ ✓ TEMPREASON Tan et al. (2023) 52.8k Wikipedia/Wikidata SC Abstractive 634-2023 ✗ ✗ TemporalAlignmentQA Zhao et al. (2024) 20k Wikipedia AG Abstractive 2000-2023 ✗ ✗ ReaLTimeQA Kasai et al. (2023) 5.1k Search CS Multiple-choice 2020-2024 ✗ ✗
@@ 3 Datasets and Evaluation Benchmarks
The development of Temporal IR/QA systems fundamentally depends on the availability of temporally grounded datasets and robust evaluation methodologies used for training, testing, and benchmarking time-aware models. We provide an overview in this section of temporal datasets, organized into three categories: temporal document collections, TQA datasets, and TIR datasets.
3.1 Temporal Document Collection
Prior work has utilized diachronic and synchronic document collections as well as annotated temporal corpora.
@@ 3.1 Temporal Document Collection
Synchronous corpora represent a coherent snapshot of the world at a specific point in time. Unlike diachronic corpora, which typically span decades or years, synchronous collections capture a temporally aligned view, sometimes in conjunction with structured knowledge bases. Wikipedia articles Vrandečić and Krötzsch (2014), for example, reflect a particular version of world knowledge at a certain time (when the dump was made) and can be linked to Wikidata timestamps. Datasets like TimeQA Chen et al. (2021), TEMPREASON Tan et al. (2023), and ComplexTempQA Gruber et al. (2024) build on Wikipedia snapshots to support temporally-scoped QA grounded in a time-specific context.
Finally, Annotated temporal corpora with explicit temporal annotations facilitate more structured forms of temporal reasoning. TimeBank Pustejovsky et al. (2003) introduced TimeML to annotate temporal expressions, events, and their temporal relations. Follow-up datasets like WikiWars Mazur and Dale (2010) and RED O’Gorman et al. (2016) extended it to historical narratives and causal relations, respectively. Such corpora constitute gold-standard resources for temporal tagging and relation extraction.
@@ 3.2 TQA Datasets
TQA datasets allow evaluating how well systems can answer questions that require temporal reasoning. They vary along multiple dimensions, including Knowledge Source, Temporal Orientation, Temporal Explicitness, and Reasoning Complexity.
Knowledge Source
TQA datasets are commonly derived from diachronic or synchronic corpora. Diachronic Corpora (also known as Primary Sources) tend to provide contemporaneous accounts written around the time when events occurred in the past. Datasets such as NewsQA Trischler et al. (2017), TDDiscourse Naik et al. (2019), TORQUE Ning et al. (2020), ArchivalQA Wang et al. (2022), TKGQA Ong et al. (2023), ChroniclingAmericaQA Piryani et al. (2024b), are curated from old news sources and can be used to evaluate models’ abilities to retrieve and reason over temporally anchored document collections. Table 1 lists all the datasets and the types of knowledge sources used to generate their questions.
In contrast, Synchronic Corpora like Wikipedia basically constitute Secondary Sources since they provide retrospective view of the past. They have been used to build datasets like TimeQA Chen et al. (2021), TEMPREASON Tan et al. (2023), TiQ Jia et al. (2024), and ComplexTempQA Gruber et al. (2024), which support fine-grained reasoning across temporally scoped, consistent knowledge bases.
Recent advancements have also seen the emergence of purely synthetic datasets designed to specifically test models on controlled and complex temporal reasoning scenarios. For example, UnSeenTimeQA Uddin et al. (2024) introduces a novel, data contamination-free benchmark that evaluates temporal reasoning independently from any pre-training knowledge.
Temporal Orientation
While most datasets focus on past events, future-oriented QA datasets remain relatively rare. Still, they are increasingly important for evaluating models’ ability to perform predictive and hypothetical reasoning. ForecastQA Jin et al. (2021) and TimeBench Chu et al. (2024) are among the few benchmarks that include questions about future events, testing models’ ability to perform timeline projections and forecast-based inference.
Question Type
Temporal questions can be broadly classified by their explicitness in referencing time. Datasets like TimeQA Chen et al. (2021), SituatedQA Zhang and Choi (2021) and TempQuestions Jia et al. (2018) contain Explicit Temporal Questions with clear temporal markers, such as "What happened in 1947?", signaling temporal intent directly.
@@
TQA tasks also vary in the depth of reasoning they require. Simple Temporal Questions typically involve direct lookups, such as identifying the date of a specific event or the state of the world at a given time. Early datasets like NewsQA Trischler et al. (2017) and TempLAMA Dhingra et al. (2022) largely belong to this category. In contrast, Complex Temporal Questions demand more intricate processing such as multi-hop reasoning, temporal filtering, or synthesizing information across events. For example, the question “What major international agreements were signed after World War I but before World War II?” necessitates multi-hop temporal reasoning and contextual comparison. Datasets like MenatQA Wei et al. (2023), TempReason Tan et al. (2023), Complex-TR Tan et al. (2024), and ComplexTempQA Gruber et al. (2024) are explicitly designed to evaluate these advanced reasoning capabilities. Others like TimeBench Chu et al. (2024) span both simple and complex reasoning levels, including tasks such as timeline construction or event duration inference. Table 1 compares various datasets for Temporal QA/IR.
3.3 TIR Datasets
While TQA datasets focus on answering time-sensitive questions, TIR datasets support tasks such as identifying time-sensitive documents, modeling temporal query intent, and ranking documents by temporal relevance or diversity. They typically pair queries with timestamped corpora and are designed to assess retrieval systems’ performance across temporal dimensions.
The Temporalia series at NTCIR-11 and NTCIR-12 Joho et al. (2014, 2016) established foundational benchmarks for TIR through two tasks: Temporal Query Intent Classification (TQIC), which categorizes queries by temporal orientation (e.g., past, recency, future, atemporal), and TIR, which ranks documents based on their temporal relevance or diversity. The tasks use the LivingKnowledge News/Blog Corpus Matthews et al. (2010), containing 3.8 million timestamped documents (2011–-2013) annotated with time expressions and named entities. Apart from Temporalia, TREC Temporal Summarization Track Diaz et al. (2015) offered datasets for a related task of real-time event summarization, testing systems’ ability to rank documents by recency and relevance as well as emphasizing temporal diversity and freshness. In parallel, the TempEval series from the SemEval workshops UzZaman et al. (2013); Verhagen et al. (2010, 2007) provided benchmark datasets for temporal information extraction such as temporal expression, event, and temporal relation, crucial for supporting TIR tasks.
4 Temporal Prediction Tasks
Temporal prediction tasks are essential for developing time-aware IR and QA systems. They focus on inferring implicit or missing temporal information from text, thereby improving the alignment between queries, documents, and events. These tasks are critical when explicit temporal metadata is sparse, noisy, or unavailable, and they support applications such as historical search, timeline construction, and temporally sensitive retrieval.
Key tasks include Event Dating, Document Dating, Focus Time Estimation, Query Time Profiling, and Event Occurrence Prediction. Traditional methods rely on statistical language models and handcrafted rules, while more recent techniques employ transformer-based encoders, temporal embeddings, and graph-based reasoning to improve generalization and robustness Yang et al. (2023); Abdallah et al. (2025); Liu and Quan (2025); Yang et al. (2024). For a detailed review of task definitions, representative techniques, and evaluation strategies, we refer the readers to Appendix C.
5 Approaches in Temporal IR/QA

### v3

# It’s High Time : A Survey of Temporal Question Answering

@@ Abstract
Time plays a critical role in how information is generated, retrieved, and interpreted. In this survey, we provide a comprehensive overview of Temporal Question Answering (TQA), a research area that focuses on answering questions involving temporal constraints or context. As the amount of time-stamped content from sources like news articles, web archives, and knowledge bases increases, systems must address challenges such as detecting temporal intent, normalizing time expressions, ordering events, and reasoning over evolving or ambiguous facts. We focus on recent advances in TQA enabled by neural architectures, especially transformer-based models and Large Language Models (LLMs), highlighting progress in temporal language modeling, retrieval-augmented generation (RAG), and temporal reasoning. We also discuss benchmark datasets and evaluation strategies designed to test temporal robustness, recency awareness, and generalization.
@@
: A Survey of Temporal Question Answering
Bhawna Piryani,♠ Abdelrahman Abdallah,♠ Jamshid Mozafari,♠ Avishek Anand,♡ Adam Jatowt♠
♠University of Innsbruck
{bhawna.piryani, abdelrahman.abdallah,
jamshid.mozafari, adam.jatowt}@uibk.ac.at
♡TU Delft
avishek.anand@tudelft.nl
@@ 1 Introduction
From analyzing centuries-old texts to answering questions about emerging developments, time shapes how we seek and interpret information. As digital content grows across time-stamped sources like news archives, social media, and knowledge bases, the ability to process and reason over temporal information has become essential Alonso et al. (2007). Temporal Question Answering (TQA) is a subfield of QA that focuses on answering questions constrained by or dependent on time, requiring systems to go beyond surface-level retrieval and processing and engage in temporal understanding and reasoning.
TQA faces distinct challenges that set it apart from standard QA tasks. These include temporal ambiguity resolution, where vague expressions like "recently" or "after the war" must be interpreted relative to the document or query context; cross-temporal reasoning, which requires understanding causal and sequential relationships between events across time; and knowledge volatility, where the continual evolution of facts renders static corpora and pre-trained models inadequate for up-to-date temporal reasoning. Temporal intent may be expressed explicitly (e.g., "post-World War II", "in 1998") or implicitly, requiring systems to infer appropriate timeframes from context. Effectively addressing these challenges demands advanced temporal reasoning, context-aware retrieval strategies, and mechanisms for dynamically adapting to changing knowledge (Berberich et al., 2010).
These challenges are illustrated in Figure 1, which presents two example questions. Q1: "At what age did Obama win the Nobel Peace Prize?" requires identifying and grounding two temporal anchors, Obama’s birth year (1961) and the year he received the Nobel Peace Prize (2009), followed by computing the interval between these two events to derive the answer: 48 years old.
Q2: "What does President Obama’s climate policy tell us about how the U.S. viewed climate change during his late years of service?" requires a more nuanced understanding of temporal context. The phrase "late years of service" requires anchoring to Obama’s presidency timeline (2009-2017). Accurate answering necessitates accessing contemporaneous policy documents from that specific period, rather than relying on retrospective analyses, which highlights how the temporal context determines which knowledge sources are relevant. As shown in the figure, the model must interpret relative expressions such as "today" or "next week" that appear in document content, anchoring them to the document’s publication date (August 4, 2015) to properly contextualize Obama’s climate policy statements within their historical moment.
@@
Research in TQA has evolved significantly, transitioning from rule-based pipelines Harabagiu and Bejan (2005) and statistical models Berberich et al. (2010); Wang et al. (2020) to modern neural systems capable of complex temporal reasoning Dhingra et al. (2022); Wang et al. (2023). Early approaches relied on handcrafted rules to normalize time expressions and resolve event sequences, but lacked scalability and struggled with diverse temporal contexts. The emergence of pre-trained language models has transformed the field by enabling robust temporal reasoning Jain et al. (2023), temporal event sequencing Lin et al. (2021), and adaptation to evolving knowledge through continual learning Han et al. (2021). These advancements paved the way for more dynamic systems that perform multi-hop temporal inference and incorporate up-to-date information through retrieval augmentation.
Despite this progress, current systems still face challenges such as handling future-oriented questions, reasoning over temporally inconsistent documents, and mitigating knowledge decay. Addressing these problems requires a deeper understanding of the field’s current capabilities and limitations.
While prior surveys111See Section A for a detailed discussion. have addressed general QA/IR techniques Kolomiyets and Moens (2011); Zhu et al. (2023, 2021) or narrow aspects of temporal processing Kobayashi and Takeda (2000); Kanhabua et al. (2015); Campos et al. (2014), a dedicated and up-to-date synthesis of TQA is missing. The most recent work to partially cover TQA appeared nearly a decade ago Campos et al. (2014), before the rise of pre-trained language models and large-scale temporal benchmarks that now define the field.
In this survey, we present a comprehensive synthesis of recent work in TQA from an unstructured text. Our scope is limited to models and datasets that operate over natural language questions and unstructured documents, excluding structured approaches based on knowledge graphs (KGs); for KG-based methods, readers may refer to Su et al. (2024a); Saxena et al. (2021). Similarly, while temporal reasoning is a component of TQA, comprehensive coverage of temporal reasoning datasets and benchmarks is beyond our scope; for extensive analysis of temporal reasoning capabilities, we refer readers to Wenzel and Jatowt (2023).
We review the evolution of TQA methods, benchmark datasets, evaluation strategies, and emerging challenges. To guide this discussion, we present a taxonomy, as illustrated in Figure 2 in the Appendix, categorizing TQA tasks, datasets, and approaches. Our goal is to provide a unified framework that synthesizes recent progress, clarifies open questions, and supports the design of next-generation time-aware systems.
@@ 2 Key Concepts
We first introduce the core concepts related to TQA. Temporal Information Retrieval (TIR) aims to retrieve documents that are not only topically relevant but also aligned with the query’s temporal intent. This intent may be explicit, such as “Olympics 2024”, or implicit, such as “latest Apple earnings”. TIR relies on different temporal signals, including document timestamps (publication dates), temporal expressions (“March 2023”), and event mentions (“2024 Olympics”) to assess a document’s temporal relevance, which indicates how well its content aligns with the query’s intended time frame Kanhabua and Nørvåg (2008); Singh et al. (2016). While TIR facilitates time-sensitive access to documents, it often lacks mechanisms for performing deeper temporal reasoning over the content itself.
@@
On the other hand, Temporal Question Answering (TQA) focuses on answering questions with temporal constraints, either explicitly stated, such as "Who won the Nobel Prize in Physics in 2020?", or implied, for instance, "What are the latest US climate policies?". Success in TQA requires understanding the question’s temporal intent and retrieving documents relevant to the corresponding time frame, or published around that time. Unlike TIR, TQA often involves more sophisticated processing, including interpreting temporal expressions, ordering events, and grounding answers in the appropriate time frame. As a result, it frequently requires multi-hop reasoning across documents and temporal signals.
@@
TQA relies on diverse temporal cues. Temporal signals include explicit temporal expressions like "March 2023", relative expressions such as "last week", implicit cues like "recently", and event-based references like "2024 Olympics". These require contextual interpretation and sometimes external temporal knowledge. Temporal metadata, such as document timestamps, indicate publication time and often serve as proxies for judging the freshness of content.
A critical yet often underutilized concept is that of document focus time, the time period a document primarily discusses, which may differ from its publication timestamp. For example, a news article published in 2013 about the 2010 Academy Awards has a publication date of 2013 but a focus time of 2010. This distinction is crucial in TQA, where users often seek information tied to specific historical or future contexts rather than content merely anchored to when it was written.
Estimating focus time is non-trivial, often requiring temporal expression normalization, burst detection, or event timestamping. When accurately modeled, it improves retrieval quality in time-sensitive settings, especially for long-range archives or temporally anchored questions. Prior work has shown that focus time estimation can enhance performance in tasks such as historical QA, event ordering, and temporally grounded search Jatowt et al. (2013); Wang et al. (2020).
@@
Dataset #Questions Knowledge Source Creation Method Answer Type Time Frame Temporal Metadata Multi-Hop NewsQA Trischler et al. (2017) 119k News CS Freeform 2007-2015 ✗ ✗ TDDiscourse Naik et al. (2019) 6.1k News CS Extractive Unspecified ✗ ✗ TORQUE Ning et al. (2020) 21k News CS Abstractive - ✗ ✗ ArchivalQA Wang et al. (2022) 532k News AG Extractive 1987-2007 ✓ ✗ TimeQA Chen et al. (2021a) 41.2K Wikipedia AG Extractive 1367-2018 ✗ ✗ TiQ Jia et al. (2024) 10K Wikipedia AG Freebase Unspecified ✗ ✗ TempQuestions Jia et al. (2018) 1.2k Freebase AG Extractive Unspecified ✗ ✓ TemporalQuestions Wang et al. (2021a) 1K News CS Extrcative 1987-2007 ✓ ✗ TempLAMA Dhingra et al. (2022) 50k News CS Extractive 2010-2020 ✓ ✗ ComplexTempQA Gruber et al. (2024) 100,228k Wikipedia AG Extractive 1987-2023 ✓ ✓ MenatQA Wei et al. (2023) 2.8k Wikipedia AG Extractive 1367-2018 ✗ ✗ PAT-Question Meem et al. (2024) 6,1k Wikipedia CS Extractive - ✗ ✓ TempTabQA Gupta et al. (2023) 11.4k Wikipedia Info box CS Abstractive - ✗ ✗ SituatedQA Zhang and Choi (2021) 12.2k Wikipedia CS - 2021 ✗ ✗ UnSeenTimeQA Uddin et al. (2024) 3.6k Synthetic AG Abstractive - ✗ ✓ ChroniclingAmericaQA Piryani et al. (2024b) 485k News AG Extractive 1800-1920 ✓ ✗ FRESHQA Vu et al. (2024) 600 Google Search CS - - ✗ ✓ COTEMPQA Su et al. (2024b) 4.7k Wikidata CS Abstractive 2023 ✗ ✓ Test of Time (ToT) Fatemi et al. (2024) 1.8k Synthetic AG Abstractive - ✗ ✓ TIMEDAIL Qin et al. (2021) 1.1k DailyDialog CS Multiple-choice - ✗ ✗ Complex-TR Tan et al. (2024) 10.8 Wikipedia+Google Search AG Multi-answer 2023 ✗ ✓ StreamingQA Liska et al. (2022) 147k News Cs Extractive 2007-2020 ✓ ✓ TRACIE Zhou et al. (2021) 5.4k Wikipedia CS abstractive 2020 ✗ ✗ ForecastQA Jin et al. (2021) 10.3k News CS Multiple-Choice 2015-2019 ✓ ✓ TEMPREASON Tan et al. (2023) 52.8k Wikipedia/Wikidata SC Abstractive 634-2023 ✗ ✗ TemporalAlignmentQA Zhao et al. (2024) 20k Wikipedia AG Abstractive 2000-2023 ✗ ✗ ReaLTimeQA Kasai et al. (2023) 5.1k Search CS Multiple-choice 2020-2024 ✗ ✗
@@ 3 Datasets and Evaluation Benchmarks
Building on the temporal representation concepts introduced in Section 2, we now examine the resources and benchmarks used to evaluate systems that address temporally grounded questions. The development of TQA systems fundamentally depends on the availability of temporally grounded datasets and robust evaluation methodologies used for training, testing, and benchmarking time-aware models. This section provides an overview of temporal datasets, organized into two categories: temporal document collections that serve as knowledge sources, and TQA datasets that enable systematic evaluation of temporal reasoning capabilities.
3.1 Temporal Document Collections
Prior work has utilized three types of temporally structured document collections, each offering distinct advantages for temporal reasoning tasks.
@@ 3.1 Temporal Document Collections
Synchronic corpora represent a coherent snapshot of the world at a specific point in time. Unlike diachronic corpora, which typically span decades or years, synchronous collections capture a temporally aligned view, sometimes in conjunction with structured knowledge bases. Wikipedia articles Vrandečić and Krötzsch (2014), for example, reflect a particular version of world knowledge at a certain time (when the dump was made) and can be linked to Wikidata timestamps. Datasets like TimeQA Chen et al. (2021a), TEMPREASON Tan et al. (2023), and ComplexTempQA Gruber et al. (2024) build on Wikipedia snapshots to support temporally-scoped QA grounded in a time-specific context.
Finally, Annotated temporal corpora with explicit temporal annotations facilitate more structured forms of temporal reasoning. TimeBank Pustejovsky et al. (2003) introduced TimeML to annotate temporal expressions, events, and their temporal relations. Follow-up datasets like WikiWars Mazur and Dale (2010) and RED O’Gorman et al. (2016) extended it to historical narratives and causal relations, respectively.
These corpora serve as gold-standard resources for temporal tagging, relation extraction, and developing models that can explicitly reason about temporal structures in text.
@@ 3.2 TQA Datasets
TQA datasets enable the evaluation of how well systems can answer questions that require temporal understanding and reasoning. They vary along multiple dimensions, including Knowledge Source, Temporal Orientation, Temporal Explicitness, and Reasoning Complexity. Table 1 provides a comprehensive comparison of major TQA datasets across these dimensions.
Knowledge Source: TQA datasets are commonly derived from diachronic or synchronic corpora. Primary Source Dataset derive from diachronic corpora, providing contemporaneous accounts written when events occurred. These datasets test models’ ability to retrieve and reason over temporally anchored document collections with authentic historical perspectives. Datasets such as NewsQA Trischler et al. (2017), TDDiscourse Naik et al. (2019), TORQUE Ning et al. (2020), ArchivalQA Wang et al. (2022), TKGQA Ong et al. (2023), ChroniclingAmericaQA Piryani et al. (2024b), are curated from historical news sources and support temporal reasoning over period-specific content.
In contrast, Synchronic Corpora like Wikipedia basically constitute Secondary Sources since they provide retrospective view of the past. They have been used to build datasets like TimeQA Chen et al. (2021a), TEMPREASON Tan et al. (2023), TiQ Jia et al. (2024), and ComplexTempQA Gruber et al. (2024), which support fine-grained reasoning across temporally scoped, consistent knowledge bases.
Temporal Orientation: While most datasets focus on past events, future-oriented QA datasets remain relatively rare. Still, they are increasingly important for evaluating models’ ability to perform predictive and hypothetical reasoning. ForecastQA Jin et al. (2021), FutureContext Mutschlechner and Jatowt (2025) and TimeBench Chu et al. (2024) are among the few benchmarks that include questions about future events, testing models’ ability to perform timeline projections and forecast-based inference.
Question Type: Temporal questions can be broadly classified by their explicitness in referencing time. Datasets like TimeQA Chen et al. (2021a), SituatedQA Zhang and Choi (2021) and TempQuestions Jia et al. (2018) contain Explicit Temporal Questions with clear temporal markers, such as "What happened in 1947?", signaling temporal intent directly.
@@
Temporal Reasoning Complexity: TQA tasks also vary in the depth of reasoning they require. Simple Temporal Questions typically involve direct lookups, such as identifying the date of a specific event or the state of the world at a given time. Early datasets like NewsQA Trischler et al. (2017) and TempLAMA Dhingra et al. (2022) largely belong to this category. In contrast, Complex Temporal Questions demand more intricate processing such as multi-hop reasoning, temporal filtering, or synthesizing information across events. For example, the question “What major international agreements were signed after World War I but before World War II?” necessitates multi-hop temporal reasoning and contextual comparison. Datasets like MenatQA Wei et al. (2023), TempReason Tan et al. (2023), Complex-TR Tan et al. (2024), and ComplexTempQA Gruber et al. (2024) are explicitly designed to evaluate these advanced reasoning capabilities. Others like TimeBench Chu et al. (2024) span both simple and complex reasoning levels, including tasks such as timeline construction or event duration inference.
Summary: TQA poses challenges that go beyond annotation quality or dataset scale. Unlike static QA tasks, the answers to temporal questions can shift over time, necessitating benchmarks that can account for answer drift and temporal volatility. Most existing datasets are static snapshots; notable exceptions include RealTimeQA Kasai et al. (2023) and FreshQA Vu et al. (2024), which incorporate periodic updates, though these updates are often resource-intensive. Alternatively, datasets like PATQA Meem et al. (2024) explore scalable, automated update mechanisms. A persistent issue is temporal ambiguity Piryani et al. (2024a), where missing or implicit time references hinder both annotation and evaluation. Structural biases also emerge: diachronic corpora tend to have more data surrounding major events and contain detailed information on past events, yet they reflect contemporary perspectives. In contrast, synchronic corpora provide broader coverage but often lack fine-grained event granularity. Annotation strategies further involve trade-offs: crowdsourced datasets are typically small but high-quality, whereas automatically generated datasets are larger but noisier. Moreover, most benchmarks fail to isolate specific reasoning skills such as duration inference or event ordering hereby limiting diagnostic evaluation.
While this section focuses on general-domain datasets, specialized domains such as medical, legal, and financial contexts present unique temporal reasoning challenges and dataset requirements, as shown in Appendix E. To support meaningful progress, robust TQA datasets must evolve over time, resolve temporal ambiguity, and facilitate fine-grained reasoning assessment.
4 Approaches in Temporal QA